In 1993, Blossom Company completed the construction of a building at a cost of $2,460,000 and first occupied it in January 1994. It was estimated that the building will have a useful life of 40 years and a salvage value of $74,400 at the end of that time.

Early in 2004, an addition to the building was constructed at a cost of $615,000. At that time, it was estimated that the remaining life of the building would be, as originally estimated, an additional 30 years, and that the addition would have a life of 30 years and a salvage value of $24,600.

In 2022, it is determined that the probable life of the building and addition will extend to the end of 2053, or 20 years beyond the original estimate.

(A) Using the straight-line method, compute the annual depreciation that would have been charged from 1994 through 2003.

Annual depreciation from 1994 through 2003

$___/yr

(B)Compute the annual depreciation that would have been charged from 2004 through 2022.

Annual depreciation from 2004 through 2021

$

/ yr

(C) Prepare the entry, if necessary, to adjust the account balances because of the revision of the estimated life in 2021.

(D) Compute the annual depreciation to be charged, beginning with 2022. (Round answer to 0 decimal places, e.g. 45,892.)

Annual depreciation expense—building

$
Answer #1
A
Cost 2460000
Less: Salvage value 74400
Depreciable cost 2385600
Useful life 40
Annual depreciation from 1994 through 2003 59640 / yr
B
Cost 2460000
Add: Additions 615000
Total cost 3075000
Less: Accumulated depreciation from 1994 through 2003 596400 =59640*10
Book value, Jan 2004 2478600
Less: Salvage value 99000 =74400+24600
Depreciable cost 2379600
Useful life 30
Annual depreciation from 2004 through 2021 79320 / yr
C
No entry is required because of the revision of the estimated life in 2021
No entry                       0
         No entry                            0
D
Book value, Jan 2004 2478600
Less: Accumulated depreciation from 2004 through 2021 1427760 =79320*18
Book value, Jan 2022 1050840
Less: Salvage value 99000 =74400+24600
Depreciable cost 951840
Useful life 32 =50-18
Annual depreciation expense—building 29745